# On integer distance sets

### v2

@@ 3.1. An explicit description of $\overline{X_{k}}$
To prove Lemma 3.2, it suffices, by standard results (e.g., combine Theorems 4 and 8 of [7, Chapter 8 §4]), to show that $Q_{m,P_{1}},\dots,Q_{m,P_{k}}$ is a Gröbner basis [^1] See Definition 5 in [7, Chaper 2 §5] for the definition of a Gröbner basis. for $I:=\langle Q_{m,P_{1}},\dots,Q_{m,P_{k}}\rangle$ with respect to some graded monomial order. Recall that a monomial order is a total order $\succ$ on $\mathcal{M}$ satisfying
@@
Finally, recall that the multivariable polynomial division algorithm (see, for example, [7, Chapter 2 §3]) says that the remainder of $Q$ under division by a set of polynomials $G=\{g_{1},\dots,g_{n}\}\subset\mathbf{C}[x,y,d_{1},\dots,d_{k}]$ is zero if and only if we can write
@@
We will also require Buchberger’s criterion (see [7, Chapter 2]), which gives a convenient way to check whether a generating set of an ideal is a Gröbner basis.
@@ 3.2. Dimension and degree of $\overline{X_{k}}$
Next, we will confirm that $\overline{X_{k}}$ is a surface of degree $2^{k}$. There are a few ways to see that $\overline{X_{k}}$ is a projective surface, but perhaps the fastest is to note that $\dim{\overline{X_{k}}}=\dim{X_{k}}$, and to use the general fact (see [7, Chapter 9 §3]) that if $J\subset\mathbf{C}[x,y,d_{1},\dots,d_{k}]$ is an ideal and $G=\{g_{1},\dots,g_{\ell}\}$ is a Gröbner basis for $J$ with respect to a graded monomial order, then $\dim{V(J)}$ equals the maximal size of a subset $A$ of the variables $\{x,y,d_{1},\dots,d_{k}\}$ such that, for all $j=1,\dots,\ell$, the leading term of $g_{j}$ contains a variable not in $A$. The following lemma is then an immediate consequence of Lemma 3.4 (by taking $A=\{x,y\}$).
@@
Now, we will show that $\overline{X_{k}}$ has degree $2^{k}$. This will follow easily from two standard facts, which we record here for later use. Both are immediate consequences of Theorem 7.7 from [18, Chapter I].
@@
By Lemmas 3.2 and 3.6, we certainly have $\deg{\overline{X_{k}}}\leq 2^{k}$. For the lower bound, consider the linear variety $H$ of codimension $2$ defined by $H=V(x,y)$. Then
@@ 3.3. Irreducibility of $\overline{X_{k}}$
Set $Y_{k}:=X_{k}\setminus X^{\prime}_{k}$ to be the set of nonsingular points of $X_{k}$. Since $Y_{k}$ is open and dense in $X_{k}$, to show that $X_{k}$ is irreducible, it suffices to show that $Y_{k}$ is irreducible. Further, since $Y_{k}$ is a smooth quasiaffine variety (as it is a smooth, open subvariety of $X_{k}$), it is irreducible if it is connected (see, for example, Remark 7.9.1 of [18, Chapter III]). Set
@@
It follows that $Y_{k}\setminus Y^{\prime}_{k}$ must be dense in $Y_{k}$, since all irreducible components of $X_{k}$ are two-dimensional (see, for example, Proposition 7.1 of [18, Chapter I]). The proof of Lemma 3.9 thus boils down to showing that $Y_{k}\setminus Y^{\prime}_{k}$ is connected. As any set that is connected in the standard (Euclidean) topology is Zariski-connected, it thus suffices to show that $Y_{k}\setminus Y^{\prime}_{k}$ is connected in the standard topology on $\mathbf{C}^{k+2}$, which we will work in for the remainder of this section. We will next show that $Y_{k}\setminus Y^{\prime}_{k}$ is, in fact, path-connected. It will then follow that $Y_{k}$ is connected, and thus Zariski-connected, as desired.
@@
Since $H\cap W$ is homeomorphic to a (real) plane minus $k$ points, for each $A\subset\{1,\dots,k\}$ there certainly exists a path $\gamma$ contained in $H\cap W$ starting and ending at $(0,0)$ with winding number $1$ around each of the points in $\bigcup_{j\in A}(H\cap\ell_{j})$ and with winding number 00 around each of the points in $\bigcup_{j\notin A}(H\cap\ell_{j})$. We will show that each of these $2^{k}$ paths lifts to a path in $Y_{k}\setminus Y_{k}^{\prime}$ that starts from $z_{0}$ and ends at a different point in $F$; this will prove the desired statement.
@@
[…]
@@
  • If $\gamma$ has winding number 00 around the point $H\cap\ell_{j}=\left(z_{j},\frac{z_{j}}{i\sqrt{m}}\right)$, then, viewing the argument of $Q_{m,P_{j}}\big{(}\gamma(t),0\big{)}$ as a continuous function of $t$, we see that the arguments at $t=0$ and at $t=1$ are the same. Therefore, $d_{j}(1)=d_{j}(0)=\sqrt{a_{j}^{2}+mb_{j}^{2}}$.
@@ 4. Degree and irreducibility of $\overline{\mathcal{C}_{k}}$
where each $R_{j,\pm}$ is a polynomial of degree at most $j$ and, for some $20\leq j\leq d-$ 2, at least one of $R_{j,+}(w)$ or $R_{j,-}(w)$ is not identically zero.
@@
for each $0\leq j\leq d$, which shows that $\deg{R_{j,\pm}}\leq j$. If $R_{j,+},R_{j,-}\equiv 0$ for all $20\leq j\leq d-$ 2, then we must have
@@
for all $\ell=2,\dots,d$ and $k=0,\dots,\ell-2$. When $\ell=d$, this implies that $i\sqrt{m}$ and $-i\sqrt{m}$ are both roots of $Q_{d}(x,1)$ of multiplicity at least $d-1$. Thus, $(x^{2}+m)^{d-1}\mid Q_{d}(x,1)$. But, $\deg{Q_{d}(x,1)}=d$ since $\deg{Q}=d$, so we must have $2(d-1)\leq d$, which is impossible when $d\geq 3$. We conclude that $R_{j,+}\not\equiv 0$ or $R_{j,-}\not\equiv 0$ for some $20\leq j\leq d-$ 2. ∎
@@
$C^{\prime}:=\{(x,y):(x,y\sqrt{m})\in C\}$
@@ 4.1. Dimension and degree of $\overline{\mathcal{C}_{k}}$
is an immediate consequence of Bézout’s inequality (Lemma 3.6), since $\deg{\overline{Q}}=\deg{Q}=\deg{C}$. Showing that $\deg{\overline{\mathcal{C}_{k}}}\geq 2^{k}$ requires essentially the same argument as in the proof of Lemma 3.8, since we assumed that $C$ (and thus $C^{\prime}$) contains the origin. Let =$\ell=\{(x,y)\in\mathbf{C}^{2}:ax+by=0\}$ be a line that is not equal to $C^{\prime}$. Since the irreducible curve $C^{\prime}$ is not $\ell$, $\ell\cap C^{\prime}$ is finite. Consider the subspace $H\subset\mathbf{P}^{k+2}$ defined by $H=V(ax+by)$. Then
@@
$H_{j}=\{[x:y:z:d_{1}:\dots:d_{k}]:[x:y:z]=w_{j}\}$
for some fixed $w_{j}\in\mathbf{P}^{2}$ (a point in the finite set ¯$\overline{\ell}\cap\overline{C^{\prime}}$). Since, for each $w\in\mathbf{P}^{2}$, the fiber
@@ 4.2. Irreducibility of $\overline{\mathcal{C}_{k}}$
First of all, it is a standard fact that the set of singular points of an algebraic curve is finite (see, for example, Theorem 5.3 of [18, Chapter I]).
@@
has the form $P^{\prime}=(x,y,(-1)^{\epsilon_{1}}d_{1},\dots,(-1)^{\epsilon_{k}}d_{k})$ for some $(\epsilon_{1},\dots,\epsilon_{k})\in\{0,1\}^{k}$. Then $J(P^{\prime})=J(P)A$ for the $(k+2)\times(k+2)$ diagonal matrix $A$ that has entries $1,1,(-1)^{\epsilon_{1}},\dots,(-1)^{\epsilon_{k}}$ along the diagonal. Since $A$ is nonsingular, $J(P^{\prime})$ must have the same rank as $J(P)$. It follows that $P^{\prime}\in\mathcal{C}_{k}^{\prime}$ as well, proving (4.3). ∎
@@
To prove (3), first note that $\overline{C^{\prime}}=V(\overline{Q})$. By Bézout’s theorem, the number of intersections, counted with multiplicity, of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$ are both $d:=\deg{Q}$. Letting $n_{1},n_{2},n_{1}^{\prime},$ and $n_{2}^{\prime}$ denote the intersection multiplicity of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ at $[a_{j}:b_{j}:1]$ and at a point at infinity and with $\overline{\ell_{j}^{\prime}}$ at $[a_{j}:b_{j}:1]$ and at a point at infinity, respectively, it thus suffices to show that $n_{1}+n_{2}<d$ or $n_{1}^{\prime}+n_{2}^{\prime}<d$.
Let us first consider the intersection multiplicities $n_{2}$ and $n_{2}^{\prime}$ of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$, respectively, at infinity when $d\geq 3$. The point at infinity lying on $\overline{\ell_{j}}$ is $[-i\sqrt{m}:1:0]$ and the point at infinity lying on $\overline{\ell_{j}^{\prime}}$ is $[i\sqrt{m}:1:0]$. On $\{[x:1:z]:x,z\in\mathbf{C}\}$, $\overline{\ell_{j}}$ and $\overline{\ell_{j}^{\prime}}$ can be parametrized as
$\{[-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\}$
@@
$\{[i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\},$
respectively. Thus, $n_{2}$ is the multiplicity of $z=0$ as a root of $Q_{+}(z)=\overline{Q}(-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z,1,z)$ and $n_{2}^{\prime}$ is the multiplicity of $z=0$ as a root of $Q_{-}(z)=\overline{Q}(i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z,1,z)$. It then follows from Lemma 4.1 and the assumption that $(a_{j},b_{j})\notin\bigcap_{l=0}^{d-2}V(\tilde{R}_{l,+},\tilde{R}_{l,-})$ that there exists a $20\leq l\leq d-$ 2 such that the coefficient of $z^{l}$ in $Q_{+}(z)$ or $Q_{-}(z)$ is nonzero. Thus, $n_{2}$ or $n_{2}^{\prime}$ is at most $d-2$ when $d\geq 3$.
@@
$-i\sqrt{m}\frac{\partial Q}{\partial x}(a_{j},b_{j})+\frac{\partial Q}{\partial y}(a_{j},b_{j})=0,$ (4.6)
@@
For the other two claims, we begin with the case that $C$, and thus $C^{\prime}$, is a line, so that $C^{\prime}=V(Q(x,y))$ where $Q(x,y)=ax+by-c$ for $a,b,c\in\mathbf{Q}$. Then no two of $\{C^{\prime},\ell_{j},\ell_{j}^{\prime}\}$ are parallel, so that $|C^{\prime}\cap\ell_{j}|=|C^{\prime}\cap\ell_{j}^{\prime}|=1$, yielding (1) and (3).
Now we deal with the case that $C$ is a circle, so that $C^{\prime}=V(Q(x,y))$ where $Q(x,y)=(x-a)^{2}+m(y-b)^{2}-r$ with $a,b,r\in\mathbf{Q}$. As in the proof of the previous lemma, we certainly have $|C^{\prime}\cap\ell_{j}|,|C^{\prime}\cap\ell_{j}^{\prime}|<\infty$ since $C^{\prime}$ is irreducible and not a line. Also as in the proof of the previous lemma, we have that the number of intersections (counted with multiplicity) of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$ are both equal to $\deg{Q}=2$. So, to conclude (3) in this case, we just need to show that it is impossible for the intersection multiplicities $n$ of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ at $[-i\sqrt{m}:1:0]$ (the point at infinity lying on $\overline{\ell_{j}}$) and $n^{\prime}$ of $\overline{C^{\prime}}$ with $\overline{\ell^{\prime}_{j}}$ at $[i\sqrt{m}:1:0]$ (the point at infinity lying on $\overline{\ell_{j}^{\prime}}$) to both equal $2$. On $\{[x:1:z]:x,z\in\mathbf{C}\}$, $\overline{\ell_{j}}$ and $\overline{\ell_{j}^{\prime}}$ can be parametrized as
$\{[-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\}$
@@
$\{[i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\},$
@@
defined as in Lemma 4.8 (3) or Lemma 4.9 (3), respectively, are all distinct. A second consequence of our observation about the topology of $C^{\prime\prime}$ is that, for any choice of $J\subset\{1,\dots,k\}$, there exists a path $\gamma_{J}$ contained in $C^{\prime\prime}$ starting and ending at $(0,0)$ with winding number $1$ around each of the points in $\{z_{j}:j\in J\}$ and with winding number 00 around each of the points in $(C^{\prime}\setminus C^{\prime\prime})\setminus\{z_{j}:j\in J\}$.
@@
$-\sqrt{Q_{m,P_{j}}\left(0,0,0\right)}=-\sqrt{a_{j}^{2}+mb_{j}^{2}},$

### v3

@@ 3.1. An explicit description of $\overline{X_{k}}$
To prove Lemma 3.2, it suffices, by standard results (e.g., combine Theorems 4 and 8 of [8, Chapter 8 §4]), to show that $Q_{m,P_{1}},\dots,Q_{m,P_{k}}$ is a Gröbner basis [^2] See Definition 5 in [8, Chaper 2 §5] for the definition of a Gröbner basis. for $I:=\langle Q_{m,P_{1}},\dots,Q_{m,P_{k}}\rangle$ with respect to some graded monomial order. Recall that a monomial order is a total order $\succ$ on $\mathcal{M}$ satisfying
@@
Finally, recall that the multivariable polynomial division algorithm (see, for example, [8, Chapter 2 §3]) says that the remainder of $Q$ under division by a set of polynomials $G=\{g_{1},\dots,g_{n}\}\subset\mathbf{C}[x,y,d_{1},\dots,d_{k}]$ is zero if and only if we can write
@@
We will also require Buchberger’s criterion (see [8, Chapter 2]), which gives a convenient way to check whether a generating set of an ideal is a Gröbner basis.
@@ 3.2. Dimension and degree of $\overline{X_{k}}$
Next, we will confirm that $\overline{X_{k}}$ is a surface of degree $2^{k}$. There are a few ways to see that $\overline{X_{k}}$ is a projective surface, but perhaps the fastest is to note that $\dim{\overline{X_{k}}}=\dim{X_{k}}$, and to use the general fact (see [8, Chapter 9 §3]) that if $J\subset\mathbf{C}[x,y,d_{1},\dots,d_{k}]$ is an ideal and $G=\{g_{1},\dots,g_{\ell}\}$ is a Gröbner basis for $J$ with respect to a graded monomial order, then $\dim{V(J)}$ equals the maximal size of a subset $A$ of the variables $\{x,y,d_{1},\dots,d_{k}\}$ such that, for all $j=1,\dots,\ell$, the leading term of $g_{j}$ contains a variable not in $A$. The following lemma is then an immediate consequence of Lemma 3.4 (by taking $A=\{x,y\}$).
@@
Now, we will show that $\overline{X_{k}}$ has degree $2^{k}$. This will follow easily from two standard facts, which we record here for later use. Both are immediate consequences of Theorem 7.7 from [21, Chapter I].
@@
By Lemmas 3.2 and 3.6, we certainly have $\deg{\overline{X_{k}}}\leq 2^{k}$. For the lower bound, consider the linear variety $H$ of codimension 22 defined by $H=V(x,y)$. Then
@@ 3.3. Irreducibility of $\overline{X_{k}}$
Set $Y_{k}:=X_{k}\setminus X^{\prime}_{k}$ to be the set of nonsingular points of $X_{k}$. Since $Y_{k}$ is open and dense in $X_{k}$, to show that $X_{k}$ is irreducible, it suffices to show that $Y_{k}$ is irreducible. Further, since $Y_{k}$ is a smooth quasiaffine variety (as it is a smooth, open subvariety of $X_{k}$), it is irreducible if it is connected (see, for example, Remark 7.9.1 of [21, Chapter III]). Set
@@
It follows that $Y_{k}\setminus Y^{\prime}_{k}$ must be dense in $Y_{k}$, since all irreducible components of $X_{k}$ are two-dimensional (see, for example, Proposition 7.1 of [21, Chapter I]). The proof of Lemma 3.9 thus boils down to showing that $Y_{k}\setminus Y^{\prime}_{k}$ is connected. As any set that is connected in the standard (Euclidean) topology is Zariski-connected, it thus suffices to show that $Y_{k}\setminus Y^{\prime}_{k}$ is connected in the standard topology on $\mathbf{C}^{k+2}$, which we will work in for the remainder of this section. We will next show that $Y_{k}\setminus Y^{\prime}_{k}$ is, in fact, path-connected. It will then follow that $Y_{k}$ is connected, and thus Zariski-connected, as desired.
@@
Since $H\cap W$ is homeomorphic to a (real) plane minus $k$ points, for each $A\subset\{1,\dots,k\}$ there certainly exists a path $\gamma$ contained in $H\cap W$ starting and ending at $(0,0)$ with winding number $1$ around each of the points in $\bigcup_{j\in A}(H\cap\ell_{j})$ and with winding number 0 around each of the points in $\bigcup_{j\notin A}(H\cap\ell_{j})$. We will show that each of these $2^{k}$ paths lifts to a path in $Y_{k}\setminus Y_{k}^{\prime}$ that starts from $z_{0}$ and ends at a different point in $F$; this will prove the desired statement.
@@
$\begin{aligned} d_{j}^{2}=Q_{m,P_{j}}\left(z,\frac{z}{i\sqrt{m}},0\right)&=2(-a_{j}+i\sqrt{m}b_{j})z+a_{j}^{2}+mb_{j}^{2}\\ &=2(-a_{j}+i\sqrt{m}b_{j})(z-z_{j})\end{aligned}$
@@
  • If $\gamma$ has winding number 0 around the point $H\cap\ell_{j}=\left(z_{j},\frac{z_{j}}{i\sqrt{m}}\right)$, then, viewing the argument of $Q_{m,P_{j}}\big{(}\gamma(t),0\big{)}$ as a continuous function of $t$, we see that the arguments at $t=0$ and at $t=1$ are the same. Therefore, $d_{j}(1)=d_{j}(0)=\sqrt{a_{j}^{2}+mb_{j}^{2}}$.
@@ 4. Degree and irreducibility of $\overline{\mathcal{C}_{k}}$
where each $R_{j,\pm}$ is a polynomial of degree at most $j$ and, for some $0\leq j\leq d-2$, at least one of $R_{j,+}(w)$ or $R_{j,-}(w)$ is not identically zero.
@@
for each $0\leq j\leq d$, which shows that $\deg{R_{j,\pm}}\leq j$. If $R_{j,+},R_{j,-}\equiv 0$ for all $0\leq j\leq d-2$, then we must have
@@
for all $\ell=2,\dots,d$ and $k=0,\dots,\ell-2$. When $\ell=d$, this implies that $i\sqrt{m}$ and $i\sqrt{m}$ are both roots of $Q_{d}(x,1)$ of multiplicity at least d1d-1. Thus, $(x^{2}+m)^{d-1}\mid Q_{d}(x,1)$. But, $\deg{Q_{d}(x,1)}=d$ since $\deg{Q}=d$, so we must have $2(d-1)\leq d$, which is impossible when $d\geq 3$. We conclude that $R_{j,+}\not\equiv 0$ or $R_{j,-}\not\equiv 0$ for some $0\leq j\leq d-2$. ∎
@@
$C^{\prime}:=\left\{(x,y):(x,y\sqrt{m})\in C\right\}$
@@ 4.1. Dimension and degree of $\overline{\mathcal{C}_{k}}$
is an immediate consequence of Bézout’s inequality (Lemma 3.6), since $\deg{\overline{Q}}=\deg{Q}=\deg{C}$. Showing that $\deg{\overline{\mathcal{C}_{k}}}\geq 2^{k}$ requires essentially the same argument as in the proof of Lemma 3.8, since we assumed that $C$ (and thus $C^{\prime}$) contains the origin. Let $\ell=\{(x,y)\in\mathbf{C}^{2}:ax+by=0\}$ be a line that is not equal to $C^{\prime}$. Since the irreducible curve $C^{\prime}$ is not $\ell$, $\ell\cap C^{\prime}$ is finite. Consider the subspace $H\subset\mathbf{P}^{k+2}$ defined by $H=V(ax+by)$. Then
@@
$H_{j}=\left\{[x:y:z:d_{1}:\dots:d_{k}]:[x:y:z]=w_{j}\right\}$
for some fixed $w_{j}\in\mathbf{P}^{2}$ (a point in the finite set $\overline{\ell}\cap\overline{C^{\prime}}$). Since, for each $w\in\mathbf{P}^{2}$, the fiber
@@ 4.2. Irreducibility of $\overline{\mathcal{C}_{k}}$
First of all, it is a standard fact that the set of singular points of an algebraic curve is finite (see, for example, Theorem 5.3 of [21, Chapter I]).
@@
has the form $P^{\prime}=(x,y,(-1)^{\epsilon_{1}}d_{1},\dots,(-1)^{\epsilon_{k}}d_{k})$ for some $(\epsilon_{1},\dots,\epsilon_{k})\in\{0,1\}^{k}$. Then $J(P^{\prime})=J(P)A$ for the $(k+2)\times(k+2)$ diagonal matrix AA that has entries $1,1,(-1)^{\epsilon_{1}},\dots,(-1)^{\epsilon_{k}}$ along the diagonal. Since $A$ is nonsingular, $J(P^{\prime})$ must have the same rank as $J(P)$. It follows that $P^{\prime}\in\mathcal{C}_{k}^{\prime}$ as well, proving (4.3). ∎
@@
To prove (3), first note that $\overline{C^{\prime}}=V(\overline{Q})$. By Bézout’s theorem, the number of intersections, counted with multiplicity, of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$ are both $d:=\deg{Q}$. Letting $n_{1},n_{2},n_{1}^{\prime},$ and $n_{2}^{\prime}$ denote the intersection multiplicity of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ at $a_{j}:b_{j}:1]$ and at a point at infinity and with $\overline{\ell_{j}^{\prime}}$ at $a_{j}:b_{j}:1]$ and at a point at infinity, respectively, it thus suffices to show that $n_{1}+n_{2}<d$ or $n_{1}^{\prime}+n_{2}^{\prime}<d$.
Let us first consider the intersection multiplicities $n_{2}$ and $n_{2}^{\prime}$ of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$, respectively, at infinity when $d\geq 3$. The point at infinity lying on $\overline{\ell_{j}}$ is $i\sqrt{m}:1:0]$ and the point at infinity lying on $\overline{\ell_{j}^{\prime}}$ is $i\sqrt{m}:1:0]$. On $\{[x:1:z]:x,z\in\mathbf{C}\}$, $\overline{\ell_{j}}$ and $\overline{\ell_{j}^{\prime}}$ can be parametrized as
$\left\{[-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\right\}$
@@
$\left\{[i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\right\},$
respectively. Thus, $n_{2}$ is the multiplicity of $z=0$ as a root of $Q_{+}(z)=\overline{Q}(-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z,1,z)$ and $n_{2}^{\prime}$ is the multiplicity of $z=0$ as a root of $Q_{-}(z)=\overline{Q}(i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z,1,z)$. It then follows from Lemma 4.1 and the assumption that $(a_{j},b_{j})\notin\bigcap_{l=0}^{d-2}V(\tilde{R}_{l,+},\tilde{R}_{l,-})$ that there exists a $0\leq l\leq d-2$ such that the coefficient of $z^{l}$ in $Q_{+}(z)$ or $Q_{-}(z)$ is nonzero. Thus, $n_{2}$ or $n_{2}^{\prime}$ is at most d2d-2 when $d\geq 3$.
@@
$i\sqrt{m}\frac{\partial Q}{\partial x}(a_{j},b_{j})+\frac{\partial Q}{\partial y}(a_{j},b_{j})=0,$ (4.6)
@@
For the other two claims, we begin with the case that $C$, and thus $C^{\prime}$, is a line, so that $C^{\prime}=V(Q(x,y))$ where Q(x,y)=ax+bycQ(x,y)=ax+by-c for $a,b,c\in\mathbf{Q}$. Then no two of $\{C^{\prime},\ell_{j},\ell_{j}^{\prime}\}$ are parallel, so that $|C^{\prime}\cap\ell_{j}|=|C^{\prime}\cap\ell_{j}^{\prime}|=1$, yielding (1) and (3).
Now we deal with the case that $C$ is a circle, so that $C^{\prime}=V(Q(x,y))$ where $Q(x,y)=(x-a)^{2}+m(y-b)^{2}-r$ with $a,b,r\in\mathbf{Q}$. As in the proof of the previous lemma, we certainly have $|C^{\prime}\cap\ell_{j}|,|C^{\prime}\cap\ell_{j}^{\prime}|<\infty$ since $C^{\prime}$ is irreducible and not a line. Also as in the proof of the previous lemma, we have that the number of intersections (counted with multiplicity) of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ and of $\overline{C^{\prime}}$ with $\overline{\ell_{j}^{\prime}}$ are both equal to $\deg{Q}=2$. So, to conclude (3) in this case, we just need to show that it is impossible for the intersection multiplicities $n$ of $\overline{C^{\prime}}$ with $\overline{\ell_{j}}$ at $i\sqrt{m}:1:0]$ (the point at infinity lying on $\overline{\ell_{j}}$) and $n^{\prime}$ of $\overline{C^{\prime}}$ with $\overline{\ell^{\prime}_{j}}$ at $i\sqrt{m}:1:0]$ (the point at infinity lying on $\overline{\ell_{j}^{\prime}}$) to both equal $2$. On $\{[x:1:z]:x,z\in\mathbf{C}\}$, $\overline{\ell_{j}}$ and $\overline{\ell_{j}^{\prime}}$ can be parametrized as
$\left\{[-i\sqrt{m}+(a_{j}+i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\right\}$
@@
$\left\{[i\sqrt{m}+(a_{j}-i\sqrt{m}b_{j})z:1:z]:z\in\mathbf{C}\right\},$
@@
defined as in Lemma 4.8 (3) or Lemma 4.9 (3), respectively, are all distinct. A second consequence of our observation about the topology of $C^{\prime\prime}$ is that, for any choice of $J\subset\{1,\dots,k\}$, there exists a path $\gamma_{J}$ contained in $C^{\prime\prime}$ starting and ending at $(0,0)$ with winding number $1$ around each of the points in $\{z_{j}:j\in J\}$ and with winding number 0 around each of the points in $(C^{\prime}\setminus C^{\prime\prime})\setminus\{z_{j}:j\in J\}$.
@@
$\sqrt{Q_{m,P_{j}}\left(0,0,0\right)}=-\sqrt{a_{j}^{2}+mb_{j}^{2}},$